OpenStudy (anonymous):

find one of the factors when x^4-3x^3-2x+6 is factored completely. a.x^3+2 b.x+3 c.x-2 d.x^3-2

hero (hero):

I could have sworn I did something exactly like this yesterday x^4-3x^3-2x+6 x^3(x-3)-2(x-3) (x-3)(x^3-2)

OpenStudy (anonymous):

\[x^4-3 x^3-2 x+6=(x-3) \left(x^3-2\right) \]